OpenStudy (anonymous):

WILL GIVE MEDAL!!! Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 6i is a zero of f(x) = x4 - 2x3 + 38x2 - 2x + 37.

OpenStudy (anonymous):

If a - ib is a zero then a+ib is a zero

OpenStudy (anonymous):

So 1+ 6i is a zero

OpenStudy (anonymous):

Here are the four roots (i, -i, 1+ 6i, 1-6i)

OpenStudy (anonymous):

Is that everytime though?

OpenStudy (anonymous):

Yes, when the polynomial has rational coefficients

OpenStudy (anonymous):

That makes sense. Thank you!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

\[ x^4 - 2 x^3 + 38 x^2 - 2 x + 37 = (1 + x^2) (37 - 2 x + x^2)=0 \] When you use the quadratic formula, then\[x= something \pm i \quad somethingelse\]